Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Intern. region
Vertex Tracker
Drift chamber
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the February 12 Background Remediation video meeting

  1. News from PEP-II
  2. The milestone of 1 Amp in the LER was achieved on February owl shift! Congratulations to the PEP-II team for this important success. The max current was 1.1 A with a lifetime of 20 minutes. The goal is 1 hour lifetime and while the lifetime increase with current is slower than expected, the availability of high currents for the remaining scrubbing week leaves lot of hopes.

    During the background commissioning meeting held on Feb 12, the various background detectors reported for the HER a factor ~2 reduction of the background and a similar reduction for the LER quadratic term compared to Dec 98. However, the LER linear term increased ~5 times for reasons not yet understood.

  3. Preparation of the collaboration meeting.
  4. A total of 3 parallel sessions is available between the two background commissioning and remediation groups. There is one plenary talk for background remediation and 1 miniplenary on Thursday. (No online people will be present on this day). (NB Two other miniplenaries have been requested for background commissioning)

    The precise topics for each session need to be defined this week.

  5. Report on event size and dead time monitoring (Gregory Dubois-Felsmann)
  6. Gregory mentioned that since several important pieces of the online software have not been installed yet, it is not possible to predict the capabilities of the system in terms of event size. There is not much action that can be done on the short term anyway. The plan is to use PRV0 MonteCarlo background frames to assess the situation, using the digi-to-TC conversion modules.

    For the dead-time monitoring, Gregory and Mike requested that a physicist be found to help them determine the minimum system that has to be put in place.

    Action item for Guy W.: Find before March 1st a physicist to work with online group on dead time monitoring issues.

  7. Report on register check back (Mike Hueffer)
  8. Mike mentioned that normally this could be handled internally in the SVT group. There could be a limitation if the SVT front ends could not interpret L1 or Read Event commands during the register check-backs. This is to be checked with David Kirkby. It should be straightforward to implement the test at the end of each run.

  9. No report was given on EPICS tools. This is thus rescheduled next week.
  10. Guy W. presented his plan for setting up a PEP-II liaison team. Comments were made that the creation of such a team should be announced in the BR plenary talks and volunteers drafted at that occasion. Discussions in a small group are needed to define scope and procedures. This will be rediscussed next week and at the next BR Steering Committee meeting on Feb 22.
  11. Carsten Hast gave some news of the SVT module. The occupancy at high LER currrent (~700 mA) is below 10 % (‘one-digit number’), the number of dead channels increased significantly during the run , up to 56%. During a 4 hour run, 1 read out incident was found.
  12. Guy W. and Tom Mattison are converging on the fast signal to be used for the protection system. The analog output of a DIRC phototube (nominal decay time 8 ns) could be used. Checks for large amplitudes are to be made.

++++Action item( GW) Contact Bob Wilson and K. Truong for information of DIRC electronics analogue chain in presence of huge signals.

9. All the talks form the Background Remediation workshop have been scanned and are now available on the BR web page.